OpenStudy (anonymous):

Use implicit differentiation to find dy/dx. 2xy-y^2=1

OpenStudy (abb0t):

\[(product. rule)-(na^{n-1})=0\]

OpenStudy (anonymous):

\[2(xy'+y)-2yy'=0\] solve for \(y'\)

OpenStudy (anonymous):

you can think of it as having \[2xf(x)-f^2(x)=0\] and being asked to solve for \(f'(x)\)

OpenStudy (anonymous):

product and chain rule gives the result

OpenStudy (anonymous):

so is the derivative just 2xy-2y?

OpenStudy (anonymous):

this answer choices are \[A. \frac{ y }{ y-x }\] \[B. \frac{ x }{ x-y }\] \[C. \frac{ x }{ y-x }\] \[D. \frac{ y }{ x-y }\]

OpenStudy (anonymous):

but how do they get one of the four choices